OpenStudy (anonymous):

tangent line equation of curve y^3 + 7y = x^3 at point (2,1) is ...

OpenStudy (welshfella):

to find the slope use implicit differentiation:- 3y^2 . dy/dx + 7 dy/dx = 3x^2 dy/dx = 3x^2 / ( 3y^2 + 7) = slope

OpenStudy (welshfella):

plug in x=2 and y=1 to evaluate the slope

OpenStudy (welshfella):

then to find the required equation use y - y1 = m(x - x1) where m = slope and (x1,x2) is the point (2,1)

OpenStudy (anonymous):

@welshfella I'm just a beginner in differentiation too I have an example in my book saying if I have an equation of y = x^3 - 3x^2 + 2x -1 for this, If I were to differentiate I would get dx/dy = 3x^2 - 6x + 2 but in this question the equation is y^3 + 7y = x^3 how do I differentiate this when there is two different variables?

OpenStudy (anonymous):

two different variables as in the y and x

OpenStudy (welshfella):

treat y as a function of x and use the chain rule if necessary so it you want to differentiate y^2 the derivative of y is dy/dx now find derivative of y^2 like you would x^2 - this gives 2y so derivative of y^2 with respect to x is 2y. dy/dx

OpenStudy (welshfella):

but maybe you haven't done the chain rule yet?

OpenStudy (anonymous):

lol yes I was about to say that I didn't really do chain rule the only thing I know so far is for instance you have 2x^2 the derivative of this will be 4x things like that..

OpenStudy (welshfella):

RIGHT WELL YOU WONT COME TO IMPLICIT DIFFERENTIATION FOR A WHILE YET

ganeshie8 (ganeshie8):

is the point really (2,1) ? it doesn't even lie on the given curve

OpenStudy (welshfella):

you use the chain rule for 'a function in a function'. for example f(x) = (x^2 + 1)^20 the x^2+1 is a function within the function f(x) so if we let u = x^2 + 1 then f(x) = u^20 then du/dx = 2x and d(fx) du = 20u^19 so d(fx) .dx = d(f(x) / du * du/dx = 2x . 20u^19 = 2x. 20(x^2 + 1)^19 = 40x(x^2 + 1)^19

OpenStudy (welshfella):

@ganeshie8 - you are right Might be a typo somewhere

OpenStudy (anonymous):

@welshfella Oh! okay Sorry just a question though... at "so d(fx) .dx = d(f(x) / du * du/dx = 2x . 20u^19" the du are getting cancelled, so I just wanna know where are you getting the 2x in 2x . 20u^19 from, if you are canceling out the du

OpenStudy (welshfella):

i'm finding that a bit difficult to explain the left side contains du/dx which we found to be 2x so we cant eliminate that we need to find d(f(x)/dx and it equals the same thing as d(fx)/dx * du / dx

OpenStudy (welshfella):

* d(fx)/du

OpenStudy (anonymous):

oh so I can't treat it like a fraction then...?

OpenStudy (welshfella):

yes you can treat it like a fraction

OpenStudy (welshfella):

we are treating it like a fraction when we are eliminating du

OpenStudy (anonymous):

yes that is exactly what I mean by treating it like a fraction when canceling du Ok so wait, the way how I understood it.. it's like multiplying 2x with 20u^19 when you said d(fx)/du * du / dx right?

OpenStudy (anonymous):

so thats why you're getting 2x . 20u^19

OpenStudy (welshfella):

yes - you got it. Better explanation than mine..

OpenStudy (anonymous):

Thank you so much Denis! :D

OpenStudy (welshfella):

yw